OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region under the curve y= (3/x+3) over the interval [0, 1] about the y-axis.

OpenStudy (zarkon):

\[2\pi\int\limits_{0}^{1}x\frac{3}{x+3}dx\]

OpenStudy (anonymous):

is there any website you can show that simplifies that a wittle further for me (:

OpenStudy (amistre64):

zarkon provided you what is sometimes called "the shell method"

OpenStudy (amistre64):

the shell method can make life easier ... in this case we could integrate by parts in one or two steps to get the results

OpenStudy (amistre64):

\begin{array}c --&--&\frac{3}{x+3}\\ +&x&3\ ln(x+3)\\ -&1&\int3\ ln(x+3) \end{array} which may or maynot be easier perhaps a usub technique?

OpenStudy (amistre64):

\[\frac{3x\ dx}{x+3}\] u = x+3 du = dx x = u-3 \[\int \frac{3(u-3)\ du}{u}\]

OpenStudy (amistre64):

this amounts to: \[3\left(\int\frac{u}{u}du-3\int\frac{1}{u}du\right)\] \[3\left(u-3\ ln(u)\right)\] \[3u-9\ ln(u)\] \[3(x+3)-9\ ln(x+3)\] \[3x+9-9\ ln(x+3) \text{ ; from 0 to 1}\]

OpenStudy (amistre64):

9+9-9 ln(4)) - ( 9 - 9 ln(3)) 18 -9 ln(4) - 9 + 9 ln(3) 9 -9 ln(4) + 9 ln(3) perhaps?

OpenStudy (amistre64):

missed it someplace; i get 6 more than wolfram

OpenStudy (amistre64):

i see it maybe; I did 3(3) instead of 3(1) :) 3(1) +9 -9ln(4) - (3(0) +9 -9ln(3)) ... thats better

OpenStudy (anonymous):

ohhhhh. Whats the final answer? lol

OpenStudy (anonymous):

hmm.. thats weird.. My choices are merely decimals of 2. something

OpenStudy (anonymous):

p.s. you sent me the solution to y= 3x/x+3 when it was x/x+3 >:| Thanks for that.

OpenStudy (zarkon):

\[x\frac{3}{x+3}=3\frac{x}{x+3}=3\frac{x+3-3}{x+3}=3\frac{x+3}{x+3}+3\frac{-3}{x+3}=3+\frac{-9}{x+3}\] then integrate

OpenStudy (anonymous):

Huh?